# Quantum geometric tensor determines the i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

### v3

@@ IV.2 Asymptotic discontinuity of QGT
Although this inequality is shown to be correct in the following subsection, its proof requires more careful analysis. This necessity arises from the asymptotic discontinuity [83, 84] of the QGT, i.e., the property that the QGT can drastically change due to a small error in conversion.
To gain a deeper understanding, let us consider the simplest case of $G=U(1)$, where its non-trivial representation is given by $e^{\mathrm{i}Ht}$ with a Hermitian operator $H$. For $G=U(1)$, the QGT is equal to the variance of the Hamiltonian $V(\psi,H)$ and its asymptotic discontinuity was pointed out in [13] and further analyzed in [23, 24, 85]. To examine the discontinuity, let $H_{N}$ be the Hamiltonian for $N$ i.i.d. copies of the system. Since the operator norm of $H_{N}$ grows linearly in $N$ and the variance is quadratic in the Hamiltonian $H_{N}$, the variance can change $O(\epsilon_{N}N^{2})$ when there is an error $\epsilon_{N}$ in states in the conversion. Therefore, the change in the variance per copy is of the order of $O(\epsilon_{N}N)$, which may not vanish as $N\to\infty$ even when $\lim_{N\to\infty}\epsilon_{N}=0$ [13, 23, 24]. Since the QGT for a generic Lie group has the same asymptotic discontinuity, a more detailed analysis is required for the proof of Eq. (34).
IV.3 Monotonicity of QGT in asymptotic conversion
To show the monotonicity of the QGT in asymptotic conversion, we investigate the asymptotic behavior of Petz’s monotone metrics near i.i.d. pure states. For this purpose, let us first review the results in [85]. For an arbitrary Hermitian operator $H$ we denote its i.i.d. extension by $H_{N}$, defined as $H_{N}\coloneqq\sum_{n=1}^{N}I^{\otimes n-1}\otimes H\otimes I^{\otimes N-n}$. For a sequence of states $\{\sigma_{N}\}_{N}$ such that $\limsup_{N\to\infty}T(\sigma_{N},\phi^{\otimes N})\leq\epsilon$ for $\epsilon>0$, and an operator monotone function $f^{(\mathrm{s})}$ satisfying $f^{(\mathrm{s})}(0)>0$ and the symmetry condition $f^{(\mathrm{s})}(t)=tf^{(\mathrm{s})}(t^{-1})$, Lemma 3 in [85] shows that under a certain regularity condition, if $\epsilon>0$ is sufficiently small, it holds
for all sufficiently large $N$, where $h$ is a real-valued function independent of $N$, satisfying $\lim_{\epsilon\to 0}h(\epsilon)=0$. The central limit theorem provides an intuitive explanation of this lemma. In the i.i.d. setting, the probability distribution of $\phi^{\otimes N}$ with respect to the eigenbasis of $H_{N}$ approaches to the normal distribution after properly redefining the random variable. It turns out that the left-hand side of Eq. (35) is approximated from below by the variance for a probability distribution modified due the the error between $\sigma_{N}$ and $\phi^{\otimes N}$. In general, modifying the distribution at values far from the mean can result in a considerable change in the variance. However, since the normal distribution has an exponentially small tail, it is impossible to significantly reduce the variance. In other words, the metric adjusted skew information approximately takes a local minimum around the i.i.d. pure states [85], expressed by Eq. (35). Note that although the definition of convergence in the state conversion in [85] is slightly different from ours, both yield the equivalent convertibility condition, as shown in Appendix B.2.
@@ IV.3 Monotonicity of QGT in asymptotic conversion
To prove the monotonicity of QGT under approximate asymptotic conversion, we show a variation of Lemma 3 in [85]. Let us first introduce several notations. When $\{\psi^{\otimes N}\}_{N}\xrightarrow{G\text{-cov.}}\{\phi^{\otimes\lfloor rN\rfloor}\}_{N}$, there exists a sequence of quantum channels $\{\mathcal{E}_{N}\}_{N}$ such that $\lim_{N\to\infty}T\left(\mathcal{E}_{N}(\psi^{\otimes N}),\phi^{\otimes\lfloor rN\rfloor}\right)=0$. For notational convenience, we define $M(N)\coloneqq\lfloor rN\rfloor$, $\sigma_{M(N)}\coloneqq\mathcal{E}_{N}(\psi^{\otimes N})$ and $\epsilon_{M(N)}\coloneqq T(\sigma_{M(N)},\phi^{\otimes M(N)})$. In the following, we omit the dependence of $M(N)$ on $N$ and simply write $M$. Since the $G$-covariance of $\mathcal{E}_{N}$ implies $\mathcal{E}_{N}(\mathcal{U}_{g}^{\otimes N}(\psi^{\otimes N}))=\mathcal{U}_{g}^{\prime\otimes M}(\sigma_{M})$, by repeating the argument presented above Eq. (32), we get $\mathcal{E}_{N}(\mathrm{i}[\psi^{\otimes N},O_{N}])=\mathrm{i}[\sigma_{M},O^{\prime}_{M}]$. Here, $O_{N}$ and $O_{M}^{\prime}$ are defined by $O_{N}\coloneqq\sum_{n=1}^{N}I^{\otimes n-1}\otimes O\otimes I^{\otimes N-n}$ and $O^{\prime}_{M}\coloneqq\sum_{n=1}^{M}I^{\otimes n-1}\otimes O^{\prime}\otimes I^{\otimes M-n}$ for $O\coloneqq\bm{\gamma}^{\dagger}\bm{X}$ and $O^{\prime}\coloneqq\bm{\gamma}^{\dagger}\bm{X}^{\prime}$. Applying Eq. (32) for $\psi^{\otimes N}$ and $\sigma_{M}$ instead of $\rho$ and $\sigma$, we find
@@
Note that the additivity of the monotone metric in the i.i.d. setting implies that the left-hand side is equal to $N\|\mathrm{i}[\psi,O]\|_{f_{q},\psi}^{2}$. To prove the monotonicity, we need to relate the right-hand side of this inequality to the QGT for $(\mathcal{U}_{g}^{\prime}(\phi))^{\otimes M}$.
@@
for all sufficiently large $N$.
Lemma 4 is proven by carefully modifying the proof of Lemma 3 in [85]. Since the proof is involved, the details are presented in Appendix B.3.
By using Lemma 4, Eq. (36) implies
$f_{q}(0)\|\mathrm{i}[\psi,O]\|_{f_{q},\psi}^{2}\geq\frac{M}{N}\bm{\gamma}^{\dagger}\mathcal{Q}^{\phi}\bm{\gamma}-\frac{M}{N}h(\epsilon)+\frac{1}{N}o\left(M\right),$ (38)
where we have used $f_{q}(0)>0$. Since $M=\lfloor rN\rfloor$, in the limit of $N\to\infty$, we get
$f_{q}(0)\|\mathrm{i}[\psi,O]\|_{f_{q},\psi}^{2}\geq r\bm{\gamma}^{\dagger}\mathcal{Q}^{\phi}\bm{\gamma}-rh(\epsilon).$ (39)
Since this inequality holds for any sufficiently small $\epsilon>0$, we get $f_{q}(0)\|\mathrm{i}[\psi,O]\|_{f_{q},\psi}^{2}\geq r\,\bm{\gamma}^{\dagger}\mathcal{Q}^{\phi}\bm{\gamma}$. In the limit of $q\to 1^{-}$, from Eq. (28), we finally get $\bm{\gamma}^{\dagger}\mathcal{Q}^{\psi}\bm{\gamma}\geq r\,\,\bm{\gamma}^{\dagger}\mathcal{Q}^{\phi}\bm{\gamma}$ for any $\bm{\gamma}\in\mathbb{C}^{\dim G}$, i.e., Eq. (34). Since $\{\psi^{\otimes N}\}_{N}\xrightarrow{G\text{-cov.}}\{\phi^{\otimes\lfloor rN\rfloor}\}_{N}$ is equivalent to $\{\mathcal{U}_{g}(\psi)^{\otimes N}\}_{N}\xrightarrow{G\text{-cov.}}\{\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN\rfloor}\}_{N}$, by repeating the argument so far for $\mathcal{U}_{g}(\psi)$ and $\mathcal{U}_{g}^{\prime}(\phi)$ instead of $\psi$ and $\phi$, we get
which completes the proof of the monotonicity of QGT in asymptotic conversion.
In summary, we have proven that $\{\psi^{\otimes N}\}_{N}\xrightarrow{G\text{-cov.}}\{\phi^{\otimes\lfloor rN\rfloor}\}_{N}$ implies $\mathcal{Q}^{\mathcal{U}_{g}(\psi)}\geq r\mathcal{Q}^{\mathcal{U}_{g}^{\prime}(\phi)}$ for all $g\in G$, which establishes the following upper bound using QGTs on the conversion rate:
Proposition 5 (Converse part).
[…]
Proposition 5 serves as the converse part of Eq. (8), i.e., the inequality showing the impossibility of achieving the rate beyond the right-hand side of Eq. (8). In the following section, we show that this rate is the optimal achievable rate.
We remark that Proposition 5 is valid even for non-compact Lie groups as long as the representation space is finite-dimensional because the proof only depends on the local properties of a state in the state space. Note also that Proposition 5 is applicable regardless of whether $\mathrm{Sym}_{G}(\psi)\subset\mathrm{Sym}_{G}(\phi)$ holds.
V QGT determines asymptotic conversion rate
Proposition 6 (Direct part).
[…]
Combining Proposition 6 with Proposition 5, our main result in Eq. (8) is proven for (non-projective) unitary representations, meaning that QGT determines the asymptotic conversion rate. As mentioned in Sec. III, this result can be extended to projective unitary representations by using the method in [59], which is detailed in Appendix A.3.
In the proof of Proposition 6, the results in quantum local asymptotic normality (QLAN) on pure-state unitary models [86, 87] play a crucial role. This can be intuitively understood through the following lemma, which provides an alternative yet equivalent characterization of the convertibility in RTA using quantum channels that are not assumed to be $G$-covariant:
Lemma 7.
Let $U$ and $U^{\prime}$ be projective unitary representations of a compact Lie group $G$ on finite-dimensional Hilbert spaces $\mathcal{H}$ and $\mathcal{H}^{\prime}$. Define $\mathcal{U}_{g}(\cdot)\coloneqq U(g)(\cdot)U(g)^{\dagger}$ and $\mathcal{U}_{g}^{\prime}(\cdot)\coloneqq U^{\prime}(g)(\cdot)U^{\prime}(g)^{\dagger}$. Let $\rho$ and $\sigma$ be arbitrary states on $\mathcal{H}$ and $\mathcal{H}^{\prime}$, respectively. For any $\epsilon\geq 0$, the following two statements are equivalent:
@@ V QGT determines asymptotic conversion rate
For $\epsilon=0$, i.e., for the conversion without error, the above is shown as Lemma 7 in [14]. Although its generalization to the case of $\epsilon>0$ is straightforward, we provide the proof in Appendix C.1 for completeness.

### v4

@@ IV.1 QGT as asymmetry monotone
However, proving this requires care because the QGT has asymptotic discontinuity [67, 68, 13, 23, 24, 69]: an asymptotically vanishing conversion error can change the QGT by $O(1)$ per copy.
This is already seen for $G=U(1)$ with representation $e^{iHt}$, where the QGT equals the energy variance [13, 23, 24, 69]. For $N$ copies, the operator norm of the Hamiltonian $H_{N}$ grows linearly with $N$. Since the variance is quadratic in $H_{N}$, a conversion error $\epsilon_{N}$ can change the variance by $O(\epsilon_{N}N^{2})$, i.e., $O(\epsilon_{N}N)$ per copy, which need not vanish even if $\epsilon_{N}\to 0$. The same issue arises for general Lie groups, hence a more detailed analysis is needed.
@@
We now prove the monotonicity of the QGT under asymptotic conversion. Suppose that $\{\psi^{\otimes N}\}_{N}\xrightarrow{G\text{-cov.}}\{\phi^{\otimes\lfloor rN\rfloor}\}_{N}$, that is, there exists a sequence of quantum channels $\{\mathcal{E}_{N}\}_{N}$ such that $\lim_{N\to\infty}T\left(\mathcal{E}_{N}(\psi^{\otimes N}),\phi^{\otimes\lfloor rN\rfloor}\right)=0$. For notational convenience, we define $M(N)\coloneqq\lfloor rN\rfloor$ and $\sigma_{M(N)}\coloneqq\mathcal{E}_{N}(\psi^{\otimes N})$. In the following, we omit the dependence of $M(N)$ on $N$ and simply write $M$. We define $O_{N}$ and $O_{M}^{\prime}$ by $O_{N}\coloneqq\sum_{n=1}^{N}I^{\otimes n-1}\otimes O\otimes I^{\otimes N-n}$ and $O^{\prime}_{M}\coloneqq\sum_{n=1}^{M}I^{\otimes n-1}\otimes O^{\prime}\otimes I^{\otimes M-n}$ with $O=\gamma^{\dagger}X$ and $O^{\prime}=\gamma^{\dagger}X^{\prime}$. Applying Eq. (30) to $\psi^{\otimes N}$ and $\sigma_{M}$ instead of $\rho$ and $\sigma$, we find
@@
By additivity of the monotone metric, the left-hand side equals $N\|\mathrm{i}[\psi,O]\|_{f_{q},\psi}^{2}$. On the other hand, the right-hand side requires careful treatment due to the asymptotic discontinuity. Nevertheless, assuming that $\lim_{M\to\infty}T\left(\sigma_{M},\phi^{\otimes M}\right)=0$, we show that there exists a real-valued function $h$, independent of $M$, with $\lim_{\epsilon\to 0}h(\epsilon)=0$, such that for any sufficiently small $\epsilon>0$,
$f_{q}(0)\|\mathrm{i}[\sigma_{M},O_{M}^{\prime}]\|_{f_{q},\sigma_{M}}^{2}\geq MV(\phi,O)-Mh(\epsilon)+o(M)$ (34)
for all sufficiently large $M$. The proof is provided in Appendix C.2.3.
From Eqs. (33) and (34) and $V(\phi,O)=\gamma^{\dagger}Q^{\phi}\gamma$, we get
$f_{q}(0)\|\mathrm{i}[\psi,O]\|_{f_{q},\psi}^{2}\geq\frac{M}{N}\gamma^{\dagger}\mathcal{Q}^{\phi}\gamma-\frac{M}{N}h(\epsilon)+\frac{o\left(M\right)}{N},$ (35)
where we used $f_{q}(0)>0$. Taking the limit $N\to\infty$ yields
$f_{q}(0)\|\mathrm{i}[\psi,O]\|_{f_{q},\psi}^{2}\geq r\gamma^{\dagger}\mathcal{Q}^{\phi}\gamma-rh(\epsilon).$ (36)
Since this inequality holds for any sufficiently small $\epsilon>0$, we get $f_{q}(0)\|\mathrm{i}[\psi,O]\|_{f_{q},\psi}^{2}\geq r\,\gamma^{\dagger}\mathcal{Q}^{\phi}\gamma$. Taking the limit of $q\to 1^{-}$ and using Eq. (28), we finally obtain Eq. (32). Since $\{\psi^{\otimes N}\}_{N}\xrightarrow{G\text{-cov.}}\{\phi^{\otimes\lfloor rN\rfloor}\}_{N}$ is equivalent to $\{\mathcal{U}_{g}(\psi)^{\otimes N}\}_{N}\xrightarrow{G\text{-cov.}}\{\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN\rfloor}\}_{N}$, by repeating the above argument for $\mathcal{U}_{g}(\psi)$ and $\mathcal{U}_{g}^{\prime}(\phi)$ instead of $\psi$ and $\phi$, we get
which completes the proof of the monotonicity of QGT in asymptotic conversion. Consequently, we obtain
Equation (38) gives the converse part of Eq. (9), i.e., it rules out any conversion rate exceeding the right-hand side of Eq. (9). In the next subsection, we show that this bound is achievable and hence optimal.
Equation (38) remains valid for non-compact Lie groups as long as the representation space is finite-dimensional, since the proof relies only on local properties of states. It also applies regardless of whether $\mathrm{Sym}_{G}(\psi)\subset\mathrm{Sym}_{G}(\phi)$ holds.
IV.2 Optimality
Here we explain key ideas to prove the optimality of the bound in Eq. (38). Essentially, we constructively show that if $r>0$ satisfies $\mathcal{Q}^{\mathcal{U}_{g}(\psi)}\geq r\,\mathcal{Q}^{\mathcal{U}^{\prime}_{g}(\phi)}$ for all $g\in G$, then any conversion rate smaller than $r$ is achievable.
The key step in proving optimality is the following lemma, which provides an alternative yet equivalent characterization of convertibility in RTA in terms of quantum channels that are not assumed to be $G$-covariant:
@@ IV.2 Optimality
Let $U$ and $U^{\prime}$ be projective unitary representations of a compact Lie group $G$ on finite-dimensional Hilbert spaces $\mathcal{H}$ and $\mathcal{H}^{\prime}$. Let $\rho$ and $\sigma$ be arbitrary states on $\mathcal{H}$ and $\mathcal{H}^{\prime}$. For any $\epsilon\geq 0$, the following are equivalent:
@@
The proof is given in Appendix D.2.1. In particular, we show that if $\mathcal{E}$ satisfies condition (i), then $\mathcal{E}^{\prime}\coloneqq\int\mathrm{d}\mu_{G}(g)\,\mathcal{U}_{g^{-1}}^{\prime}\circ\mathcal{E}\circ\mathcal{U}_{g}$ is a $G$-covariant channel and satisfies condition (ii), where $\mu_{G}$ denotes the normalized Haar measure on $G$.